OpenStudy (anonymous):

Solve: 5(x + 8) = 7(x − 2) 17 −17 27 −27

OpenStudy (anonymous):

5x+40=7x-14 54=2x 27=x

OpenStudy (anonymous):

Solve: 2(x + 1) = −4 −3 3 −2.5 −1 can you explain it to me?